1. At STP, 1 mole N2 gas occupies ____________ L, and this accounts for _____________ x 1023molecules and 28.0 ______________ of nitrogen.

2. Calculate the Kelvin temperature to which 10.0 L of a gas at 27 °C would have to be heated to change the volume to 12.0 L. Units of Kelvin are assumed here for the answer you enter so just put the number value.

3. If a plastic container of food at 1.0°C and 750. mmHg is heated in a microwave to 80. °C, what is the new pressure inside the container? mmHg are the units of the answer and you can just enter the number value for the pressure.

4. How many moles are contained in 2.0 L of N2 at standard temperature and pressure? You need just enter the number value and the units are moles of nitrogen gas.
Answer #1

1. At STP , 1 mol of every gas occupies 22.4L volume.

As at STP , temperature= 273k(0°C) and pressure= 1atm

R = 0.0821 L atm/mol K

So as , PV=nRT , V = nRT/P = 1×0.0821×273 / 1 = 22.4 L

So at STP , 1 mole N2 gas occupies 22.4L .....and accounts for 6.022 × 10^23 molecules and 28 grams of nitrogen .

1 mole of every substance is equal to its molar mass and have fix no. Of molecules in 1 mole of every substance i.e. 6.022×10^23 .